OpenStudy (anonymous):

Will FAN AND MEDAL! Please give answer and steps of how to solve it : ) A $1,600.00 principal earn 7% interest, compounded semiannually twice per year. After 33 years, what is the balance of the account.

OpenStudy (jdoe0001):

hmm actually, is 1,600 =) thus \(\qquad \textit{Compound Interest Earned Amount}\\ A=P\left(1+\frac{r}{n}\right)^{nt} \qquad \begin{cases} A=\textit{current amount}\\ P=\textit{original amount deposited}\to &\$1,600\\ r=rate\to 7\%\to \frac{7}{100}\to &0.07\\ n=\textit{times it compounds per year}\to &2\\ t=years\to &33 \end{cases}\)

OpenStudy (jdoe0001):

current amount = balance

OpenStudy (anonymous):

Let me work it out and tell me if I'm right please :) @jdoe0001

OpenStudy (jdoe0001):

well... ok but it's rather simple, just plug in the values in the compound interest formula :)

OpenStudy (anonymous):

4,979.11!! :D

OpenStudy (anonymous):

Thank you for your help!!

OpenStudy (jdoe0001):

yw